# On a counterexample to Mordell’s Pellian equation conjecture: a non-computer based approach

### v1

@@ 2. Preliminaries
${\rm v}_{P_{5}}\left(z\mathcal{O}_{K}\right)=b_{1}+b_{8}+b_{12}+b_{16}+b_{23}+b_{25}+b_{40}+b_{48}=d_{5}$,
${\rm v}_{\overline{P}_{5}}\left(z\mathcal{O}_{K}\right)=b_{7}+b_{15}+b_{30}+b_{35}+b_{36}+b_{37}+b_{52}+b_{53}=d_{5}$, ${\rm v}_{P_{6}}\left(z\mathcal{O}_{K}\right)=b_{9}+b_{11}+b_{29}+b_{32}+b_{37}=d_{6}$,
${\rm v}_{\overline{P}_{6}}\left(z\mathcal{O}_{K}\right)=b_{1}+b_{31}+b_{33}+b_{38}+b_{41}+2b_{47}+b_{49}+b_{51}+b_{53}=d_{6}$, ${\rm v}_{P_{7}}\left(z\mathcal{O}_{K}\right)=b_{4}+b_{29}+b_{39}+2b_{50}+b_{56}=d_{7}$,
${\rm v}_{\overline{P}_{7}}\left(z\mathcal{O}_{K}\right)=b_{1}+b_{2}+b_{19}+b_{34}+b_{43}+b_{54}=d_{7}$, ${\rm v}_{P_{8}}\left(z\mathcal{O}_{K}\right)=b_{15}+b_{19}+b_{22}=d_{8}$,
${\rm v}_{\overline{P}_{8}}\left(z\mathcal{O}_{K}\right)=b_{24}+b_{33}+b_{50}=d_{8}$, ${\rm v}_{P_{9}}\left(z\mathcal{O}_{K}\right)=b_{6}+b_{14}+b_{25}+b_{29}+b_{50}=d_{9}$,
${\rm v}_{\overline{P}_{9}}\left(z\mathcal{O}_{K}\right)=b_{2}+b_{40}+b_{44}+b_{53}+b_{55}=d_{9}$, ${\rm v}_{P_{10}}\left(z\mathcal{O}_{K}\right)=b_{29}+b_{38}+b_{45}=d_{10}$,
${\rm v}_{\overline{P}_{10}}\left(z\mathcal{O}_{K}\right)=b_{8}+b_{13}+2b_{34}=d_{10}$, ${\rm v}_{P_{11}}\left(z\mathcal{O}_{K}\right)=b_{26}+b_{35}+b_{42}=d_{11}$, ${\rm v}_{\overline{P}_{11}}\left(z\mathcal{O}_{K}\right)=b_{3}+b_{54}=d_{11}$,
${\rm v}_{P_{12}}\left(z\mathcal{O}_{K}\right)=b_{35}+b_{37}+b_{54}=d_{12}$, ${\rm v}_{\overline{P}_{12}}\left(z\mathcal{O}_{K}\right)=b_{4}+b_{5}+b_{7}+b_{20}+b_{25}+b_{26}=d_{12}$,
${\rm v}_{P_{13}}\left(z\mathcal{O}_{K}\right)=b_{23}+b_{29}+b_{32}+b_{42}=d_{13}$, ${\rm v}_{\overline{P}_{13}}\left(z\mathcal{O}_{K}\right)=b_{19}+b_{53}=d_{13}$, ${\rm v}_{P_{14}}\left(z\mathcal{O}_{K}\right)=b_{6}+b_{10}+b_{39}=d_{14}$,
${\rm v}_{\overline{P}_{14}}\left(z\mathcal{O}_{K}\right)=b_{54}=d_{14}$, ${\rm v}_{P_{15}}\left(z\mathcal{O}_{K}\right)=b_{13}+b_{28}+b_{34}+b_{43}+b_{46}=d_{15}$, ${\rm v}_{\overline{P}_{15}}\left(z\mathcal{O}_{K}\right)=b_{3}+2b_{19}+2b_{38}=d_{15}$,
${\rm v}_{P_{16}}\left(z\mathcal{O}_{K}\right)=b_{29}=d_{16}$, ${\rm v}_{\overline{P}_{16}}\left(z\mathcal{O}_{K}\right)=b_{2}+b_{31}+b_{51}=d_{16}$, ${\rm v}_{P_{17}}\left(z\mathcal{O}_{K}\right)=b_{12}+b_{55}=d_{17}$,
${\rm v}_{\overline{P}_{17}}\left(z\mathcal{O}_{K}\right)=b_{30}+b_{31}+b_{56}=d_{17}$, ${\rm v}_{P_{18}}\left(z\mathcal{O}_{K}\right)=b_{16}+b_{25}+b_{27}=d_{18}$,
${\rm v}_{\overline{P}_{18}}\left(z\mathcal{O}_{K}\right)=b_{21}+b_{23}+b_{39}+b_{45}=d_{18}$, ${\rm v}_{P_{19}}\left(z\mathcal{O}_{K}\right)=b_{26}+b_{29}=d_{19}$, ${\rm v}_{\overline{P}_{19}}\left(z\mathcal{O}_{K}\right)=b_{9}+b_{21}+b_{33}=d_{19}$,
${\rm v}_{P_{20}}\left(z\mathcal{O}_{K}\right)=b_{39}=d_{20}$, ${\rm v}_{\overline{P}_{20}}\left(z\mathcal{O}_{K}\right)=2b_{22}=d_{20}$, ${\rm v}_{P_{21}}\left(z\mathcal{O}_{K}\right)=b_{31}+b_{44}+b_{46}=d_{21}$,
${\rm v}_{\overline{P}_{21}}\left(z\mathcal{O}_{K}\right)=b_{13}+2b_{18}+b_{41}=d_{21}$, ${\rm v}_{P_{22}}\left(z\mathcal{O}_{K}\right)=2b_{23}=d_{22}$, ${\rm v}_{\overline{P}_{22}}\left(z\mathcal{O}_{K}\right)=b_{28}+b_{52}+b_{57}=d_{22}$,
${\rm v}_{P_{23}}\left(z\mathcal{O}_{K}\right)=b_{14}+b_{16}+b_{37}+b_{44}=d_{23}$, ${\rm v}_{\overline{P}_{23}}\left(z\mathcal{O}_{K}\right)=b_{6}+b_{28}+b_{39}+b_{42}=d_{23}$, ${\rm v}_{P_{24}}\left(z\mathcal{O}_{K}\right)=b_{30}=d_{24}$,
${\rm v}_{\overline{P}_{24}}\left(z\mathcal{O}_{K}\right)=b_{50}=d_{24}$, ${\rm v}_{P_{25}}\left(z\mathcal{O}_{K}\right)=b_{13}+b_{28}+b_{40}=d_{25}$, ${\rm v}_{\overline{P}_{25}}\left(z\mathcal{O}_{K}\right)=b_{11}+b_{52}=d_{25}$,
${\rm v}_{P_{26}}\left(z\mathcal{O}_{K}\right)=b_{32}+b_{36}=d_{26}$, ${\rm v}_{\overline{P}_{26}}\left(z\mathcal{O}_{K}\right)=b_{20}+b_{24}=d_{26}$, ${\rm v}_{P_{27}}\left(z\mathcal{O}_{K}\right)=b_{46}=d_{27}$,
${\rm v}_{\overline{P}_{27}}\left(z\mathcal{O}_{K}\right)=b_{2}+b_{49}=d_{27}$, ${\rm v}_{P_{28}}\left(z\mathcal{O}_{K}\right)=b_{22}+b_{36}+b_{55}=d_{28}$, ${\rm v}_{\overline{P}_{28}}\left(z\mathcal{O}_{K}\right)=b_{12}+b_{45}+b_{48}=d_{28}$,
${\rm v}_{P_{29}}\left(z\mathcal{O}_{K}\right)=b_{21}+b_{27}=d_{29}$, ${\rm v}_{\overline{P}_{29}}\left(z\mathcal{O}_{K}\right)=b_{36}=d_{29}$, ${\rm v}_{P_{30}}\left(z\mathcal{O}_{K}\right)=b_{25}+b_{30}+b_{49}=d_{30}$,
${\rm v}_{\overline{P}_{30}}\left(z\mathcal{O}_{K}\right)=b_{14}+b_{17}+b_{34}+b_{41}+b_{47}=d_{30}$, ${\rm v}_{P_{31}}\left(z\mathcal{O}_{K}\right)=b_{7}+b_{17}=d_{31}$, ${\rm v}_{\overline{P}_{31}}\left(z\mathcal{O}_{K}\right)=b_{45}+b_{55}=d_{31}$,
${\rm v}_{P_{32}}\left(z\mathcal{O}_{K}\right)=b_{43}+b_{45}=d_{32}$, ${\rm v}_{\overline{P}_{32}}\left(z\mathcal{O}_{K}\right)=b_{36}=d_{32}$, ${\rm v}_{P_{33}}\left(z\mathcal{O}_{K}\right)=b_{18}+b_{27}+b_{44}=d_{33}$,
${\rm v}_{\overline{P}_{33}}\left(z\mathcal{O}_{K}\right)=b_{16}+b_{24}=d_{33}$, ${\rm v}_{P_{34}}\left(z\mathcal{O}_{K}\right)=b_{33}+b_{44}=d_{34}$, ${\rm v}_{\overline{P}_{34}}\left(z\mathcal{O}_{K}\right)=b_{15}+b_{17}=d_{34}$,
${\rm v}_{P_{35}}\left(z\mathcal{O}_{K}\right)=b_{6}=d_{35}$, ${\rm v}_{\overline{P}_{35}}\left(z\mathcal{O}_{K}\right)=b_{23}=d_{35}$, ${\rm v}_{P_{36}}\left(z\mathcal{O}_{K}\right)=b_{26}+2b_{31}=d_{36}$,
${\rm v}_{\overline{P}_{36}}\left(z\mathcal{O}_{K}\right)=b_{11}+b_{34}=d_{36}$, ${\rm v}_{P_{37}}\left(z\mathcal{O}_{K}\right)=b_{4}+b_{10}+b_{15}+b_{18}+b_{19}=d_{37}$, ${\rm v}_{\overline{P}_{37}}\left(z\mathcal{O}_{K}\right)=b_{11}=d_{37}$,
${\rm v}_{P_{38}}\left(z\mathcal{O}_{K}\right)=b_{43}=d_{38}$, ${\rm v}_{\overline{P}_{38}}\left(z\mathcal{O}_{K}\right)=b_{7}+b_{21}=d_{38}$, ${\rm v}_{P_{39}}\left(z\mathcal{O}_{K}\right)=b_{5}+b_{9}+b_{35}+b_{42}+b_{46}+b_{48}+b_{55}=d_{39}$,
${\rm v}_{\overline{P}_{39}}\left(z\mathcal{O}_{K}\right)=b_{3}+b_{38}+b_{53}=d_{39}$, ${\rm v}_{P_{40}}\left(z\mathcal{O}_{K}\right)=b_{4}+b_{47}+b_{57}=d_{40}$, ${\rm v}_{\overline{P}_{40}}\left(z\mathcal{O}_{K}\right)=b_{38}=d_{40}$,
${\rm v}_{P_{41}}\left(z\mathcal{O}_{K}\right)=b_{20}=d_{41}$, ${\rm v}_{\overline{P}_{41}}\left(z\mathcal{O}_{K}\right)=b_{4}=d_{41}$, ${\rm v}_{P_{42}}\left(z\mathcal{O}_{K}\right)=b_{52}=d_{42}$, ${\rm v}_{\overline{P}_{42}}\left(z\mathcal{O}_{K}\right)=b_{30}=d_{42}$,
${\rm v}_{P_{43}}\left(z\mathcal{O}_{K}\right)=b_{1}=d_{43}$, ${\rm v}_{\overline{P}_{43}}\left(z\mathcal{O}_{K}\right)=b_{16}=d_{43}$, ${\rm v}_{P_{44}}\left(z\mathcal{O}_{K}\right)=b_{13}+b_{51}=d_{44}$,
${\rm v}_{\overline{P}_{44}}\left(z\mathcal{O}_{K}\right)=b_{12}=d_{44}$, ${\rm v}_{P_{45}}\left(z\mathcal{O}_{K}\right)=b_{51}=d_{45}$, ${\rm v}_{\overline{P}_{45}}\left(z\mathcal{O}_{K}\right)=b_{18}+b_{40}=d_{45}$,
${\rm v}_{P_{46}}\left(z\mathcal{O}_{K}\right)=b_{7}+b_{10}+b_{32}+b_{33}+b_{35}+b_{49}+b_{52}=d_{46}$, ${\rm v}_{\overline{P}_{46}}\left(z\mathcal{O}_{K}\right)=b_{2}+b_{9}+b_{42}=d_{46}$,
${\rm v}_{P_{47}}\left(z\mathcal{O}_{K}\right)=b_{6}+b_{50}+b_{54}=d_{47}$, ${\rm v}_{\overline{P}_{47}}\left(z\mathcal{O}_{K}\right)=b_{43}+2b_{56}=d_{47}$, ${\rm v}_{P_{48}}\left(z\mathcal{O}_{K}\right)=b_{9}=d_{48}$,
${\rm v}_{\overline{P}_{48}}\left(z\mathcal{O}_{K}\right)=b_{5}+b_{12}=d_{48}$, ${\rm v}_{P_{49}}\left(z\mathcal{O}_{K}\right)=b_{27}+b_{41}=d_{49}$, ${\rm v}_{\overline{P}_{49}}\left(z\mathcal{O}_{K}\right)=b_{14}+b_{40}+b_{48}=d_{49}$,
${\rm v}_{P_{50}}\left(z\mathcal{O}_{K}\right)=b_{57}=d_{50}$, ${\rm v}_{\overline{P}_{50}}\left(z\mathcal{O}_{K}\right)=b_{24}+b_{49}=d_{50}$, ${\rm v}_{P_{51}}\left(z\mathcal{O}_{K}\right)=b_{5}+b_{22}=d_{51}$,
${\rm v}_{\overline{P}_{51}}\left(z\mathcal{O}_{K}\right)=b_{41}=d_{51}$, ${\rm v}_{P_{52}}\left(z\mathcal{O}_{K}\right)=b_{3}+b_{17}+b_{21}+b_{32}+b_{48}=d_{52}$, ${\rm v}_{\overline{P}_{52}}\left(z\mathcal{O}_{K}\right)=b_{8}+b_{10}=d_{52}$,
${\rm v}_{P_{53}}\left(z\mathcal{O}_{K}\right)=b_{3}+b_{11}=d_{53}$, ${\rm v}_{\overline{P}_{53}}\left(z\mathcal{O}_{K}\right)=b_{1}+b_{14}+b_{51}+b_{57}=d_{53}$, ${\rm v}_{P_{54}}\left(z\mathcal{O}_{K}\right)=b_{8}+b_{39}=d_{54}$,
${\rm v}_{\overline{P}_{54}}\left(z\mathcal{O}_{K}\right)=b_{25}=d_{54}$, ${\rm v}_{P_{55}}\left(z\mathcal{O}_{K}\right)=b_{15}+b_{28}=d_{55}$, ${\rm v}_{\overline{P}_{55}}\left(z\mathcal{O}_{K}\right)=b_{10}+b_{35}=d_{55}$,
${\rm v}_{P_{56}}\left(z\mathcal{O}_{K}\right)=b_{37}=d_{56}$, ${\rm v}_{\overline{P}_{56}}\left(z\mathcal{O}_{K}\right)=b_{26}+b_{47}=d_{56}$.

### v2

@@ 2. Preliminaries
${\rm v}_{P_{5}}\left(z\mathcal{O}_{K}\right)=b_{1}+b_{8}+b_{12}+b_{16}+b_{23}+b_{25}+b_{40}+b_{48}=d_{5}$, ${\rm v}_{\overline{P}_{5}}\left(z\mathcal{O}_{K}\right)=b_{7}+b_{15}+b_{30}+b_{35}+b_{36}+b_{37}+b_{52}+b_{53}=d_{5}$,
${\rm v}_{P_{6}}\left(z\mathcal{O}_{K}\right)=b_{9}+b_{11}+b_{29}+b_{32}+b_{37}=d_{6}$, ${\rm v}_{\overline{P}_{6}}\left(z\mathcal{O}_{K}\right)=b_{1}+b_{31}+b_{33}+b_{38}+b_{41}+2b_{47}+b_{49}+b_{51}+b_{53}=d_{6}$,
${\rm v}_{P_{7}}\left(z\mathcal{O}_{K}\right)=b_{4}+b_{29}+b_{39}+2b_{50}+b_{56}=d_{7}$, ${\rm v}_{\overline{P}_{7}}\left(z\mathcal{O}_{K}\right)=b_{1}+b_{2}+b_{19}+b_{34}+b_{43}+b_{54}=d_{7}$, ${\rm v}_{P_{8}}\left(z\mathcal{O}_{K}\right)=b_{15}+b_{19}+b_{22}=d_{8}$,
${\rm v}_{\overline{P}_{8}}\left(z\mathcal{O}_{K}\right)=b_{24}+b_{33}+b_{50}=d_{8}$, ${\rm v}_{P_{9}}\left(z\mathcal{O}_{K}\right)=b_{6}+b_{14}+b_{25}+b_{29}+b_{50}=d_{9}$, ${\rm v}_{\overline{P}_{9}}\left(z\mathcal{O}_{K}\right)=b_{2}+b_{40}+b_{44}+b_{53}+b_{55}=d_{9}$,
${\rm v}_{P_{10}}\left(z\mathcal{O}_{K}\right)=b_{29}+b_{38}+b_{45}=d_{10}$, ${\rm v}_{\overline{P}_{10}}\left(z\mathcal{O}_{K}\right)=b_{8}+b_{13}+2b_{34}=d_{10}$, ${\rm v}_{P_{11}}\left(z\mathcal{O}_{K}\right)=b_{26}+b_{35}+b_{42}=d_{11}$, ${\rm v}_{\overline{P}_{11}}\left(z\mathcal{O}_{K}\right)=b_{3}+b_{54}=d_{11}$,
${\rm v}_{P_{12}}\left(z\mathcal{O}_{K}\right)=b_{35}+b_{37}+b_{54}=d_{12}$, ${\rm v}_{\overline{P}_{12}}\left(z\mathcal{O}_{K}\right)=b_{4}+b_{5}+b_{7}+b_{20}+b_{25}+b_{26}=d_{12}$, ${\rm v}_{P_{13}}\left(z\mathcal{O}_{K}\right)=b_{23}+b_{29}+b_{32}+b_{42}=d_{13}$,
${\rm v}_{\overline{P}_{13}}\left(z\mathcal{O}_{K}\right)=b_{19}+b_{53}=d_{13}$, ${\rm v}_{P_{14}}\left(z\mathcal{O}_{K}\right)=b_{6}+b_{10}+b_{39}=d_{14}$, ${\rm v}_{\overline{P}_{14}}\left(z\mathcal{O}_{K}\right)=b_{54}=d_{14}$, ${\rm v}_{P_{15}}\left(z\mathcal{O}_{K}\right)=b_{13}+b_{28}+b_{34}+b_{43}+b_{46}=d_{15}$,
${\rm v}_{\overline{P}_{15}}\left(z\mathcal{O}_{K}\right)=b_{3}+2b_{19}+2b_{38}=d_{15}$, ${\rm v}_{P_{16}}\left(z\mathcal{O}_{K}\right)=b_{29}=d_{16}$, ${\rm v}_{\overline{P}_{16}}\left(z\mathcal{O}_{K}\right)=b_{2}+b_{31}+b_{51}=d_{16}$, ${\rm v}_{P_{17}}\left(z\mathcal{O}_{K}\right)=b_{12}+b_{55}=d_{17}$,
${\rm v}_{\overline{P}_{17}}\left(z\mathcal{O}_{K}\right)=b_{30}+b_{31}+b_{56}=d_{17}$, ${\rm v}_{P_{18}}\left(z\mathcal{O}_{K}\right)=b_{16}+b_{25}+b_{27}=d_{18}$, ${\rm v}_{\overline{P}_{18}}\left(z\mathcal{O}_{K}\right)=b_{21}+b_{23}+b_{39}+b_{45}=d_{18}$,
${\rm v}_{P_{19}}\left(z\mathcal{O}_{K}\right)=b_{26}+b_{29}=d_{19}$, ${\rm v}_{\overline{P}_{19}}\left(z\mathcal{O}_{K}\right)=b_{9}+b_{21}+b_{33}=d_{19}$, ${\rm v}_{P_{20}}\left(z\mathcal{O}_{K}\right)=b_{39}=d_{20}$, ${\rm v}_{\overline{P}_{20}}\left(z\mathcal{O}_{K}\right)=2b_{22}=d_{20}$,
${\rm v}_{P_{21}}\left(z\mathcal{O}_{K}\right)=b_{31}+b_{44}+b_{46}=d_{21}$, ${\rm v}_{\overline{P}_{21}}\left(z\mathcal{O}_{K}\right)=b_{13}+2b_{18}+b_{41}=d_{21}$, ${\rm v}_{P_{22}}\left(z\mathcal{O}_{K}\right)=2b_{23}=d_{22}$, ${\rm v}_{\overline{P}_{22}}\left(z\mathcal{O}_{K}\right)=b_{28}+b_{52}+b_{57}=d_{22}$,
${\rm v}_{P_{23}}\left(z\mathcal{O}_{K}\right)=b_{14}+b_{16}+b_{37}+b_{44}=d_{23}$, ${\rm v}_{\overline{P}_{23}}\left(z\mathcal{O}_{K}\right)=b_{6}+b_{28}+b_{39}+b_{42}=d_{23}$, ${\rm v}_{P_{24}}\left(z\mathcal{O}_{K}\right)=b_{30}=d_{24}$, ${\rm v}_{\overline{P}_{24}}\left(z\mathcal{O}_{K}\right)=b_{50}=d_{24}$,
${\rm v}_{P_{25}}\left(z\mathcal{O}_{K}\right)=b_{13}+b_{28}+b_{40}=d_{25}$, ${\rm v}_{\overline{P}_{25}}\left(z\mathcal{O}_{K}\right)=b_{11}+b_{52}=d_{25}$, ${\rm v}_{P_{26}}\left(z\mathcal{O}_{K}\right)=b_{32}+b_{36}=d_{26}$, ${\rm v}_{\overline{P}_{26}}\left(z\mathcal{O}_{K}\right)=b_{20}+b_{24}=d_{26}$,
${\rm v}_{P_{27}}\left(z\mathcal{O}_{K}\right)=b_{46}=d_{27}$, ${\rm v}_{\overline{P}_{27}}\left(z\mathcal{O}_{K}\right)=b_{2}+b_{49}=d_{27}$, ${\rm v}_{P_{28}}\left(z\mathcal{O}_{K}\right)=b_{22}+b_{36}+b_{55}=d_{28}$, ${\rm v}_{\overline{P}_{28}}\left(z\mathcal{O}_{K}\right)=b_{12}+b_{45}+b_{48}=d_{28}$,
${\rm v}_{P_{29}}\left(z\mathcal{O}_{K}\right)=b_{21}+b_{27}=d_{29}$, ${\rm v}_{\overline{P}_{29}}\left(z\mathcal{O}_{K}\right)=b_{36}=d_{29}$, ${\rm v}_{P_{30}}\left(z\mathcal{O}_{K}\right)=b_{25}+b_{30}+b_{49}=d_{30}$, ${\rm v}_{\overline{P}_{30}}\left(z\mathcal{O}_{K}\right)=b_{14}+b_{17}+b_{34}+b_{41}+b_{47}=d_{30}$,
${\rm v}_{P_{31}}\left(z\mathcal{O}_{K}\right)=b_{7}+b_{17}=d_{31}$, ${\rm v}_{\overline{P}_{31}}\left(z\mathcal{O}_{K}\right)=b_{45}+b_{55}=d_{31}$, ${\rm v}_{P_{32}}\left(z\mathcal{O}_{K}\right)=b_{43}+b_{45}=d_{32}$, ${\rm v}_{\overline{P}_{32}}\left(z\mathcal{O}_{K}\right)=b_{36}=d_{32}$,
${\rm v}_{P_{33}}\left(z\mathcal{O}_{K}\right)=b_{18}+b_{27}+b_{44}=d_{33}$, ${\rm v}_{\overline{P}_{33}}\left(z\mathcal{O}_{K}\right)=b_{16}+b_{24}=d_{33}$, ${\rm v}_{P_{34}}\left(z\mathcal{O}_{K}\right)=b_{33}+b_{44}=d_{34}$, ${\rm v}_{\overline{P}_{34}}\left(z\mathcal{O}_{K}\right)=b_{15}+b_{17}=d_{34}$,
${\rm v}_{P_{35}}\left(z\mathcal{O}_{K}\right)=b_{6}=d_{35}$, ${\rm v}_{\overline{P}_{35}}\left(z\mathcal{O}_{K}\right)=b_{23}=d_{35}$, ${\rm v}_{P_{36}}\left(z\mathcal{O}_{K}\right)=b_{26}+2b_{31}=d_{36}$, ${\rm v}_{\overline{P}_{36}}\left(z\mathcal{O}_{K}\right)=b_{11}+b_{34}=d_{36}$,
${\rm v}_{P_{37}}\left(z\mathcal{O}_{K}\right)=b_{4}+b_{10}+b_{15}+b_{18}+b_{19}=d_{37}$, ${\rm v}_{\overline{P}_{37}}\left(z\mathcal{O}_{K}\right)=b_{11}=d_{37}$, ${\rm v}_{P_{38}}\left(z\mathcal{O}_{K}\right)=b_{43}=d_{38}$, ${\rm v}_{\overline{P}_{38}}\left(z\mathcal{O}_{K}\right)=b_{7}+b_{21}=d_{38}$,
${\rm v}_{P_{39}}\left(z\mathcal{O}_{K}\right)=b_{5}+b_{9}+b_{35}+b_{42}+b_{46}+b_{48}+b_{55}=d_{39}$, ${\rm v}_{\overline{P}_{39}}\left(z\mathcal{O}_{K}\right)=b_{3}+b_{38}+b_{53}=d_{39}$, ${\rm v}_{P_{40}}\left(z\mathcal{O}_{K}\right)=b_{4}+b_{47}+b_{57}=d_{40}$,
${\rm v}_{\overline{P}_{40}}\left(z\mathcal{O}_{K}\right)=b_{38}=d_{40}$, ${\rm v}_{P_{41}}\left(z\mathcal{O}_{K}\right)=b_{20}=d_{41}$, ${\rm v}_{\overline{P}_{41}}\left(z\mathcal{O}_{K}\right)=b_{4}=d_{41}$, ${\rm v}_{P_{42}}\left(z\mathcal{O}_{K}\right)=b_{52}=d_{42}$, ${\rm v}_{\overline{P}_{42}}\left(z\mathcal{O}_{K}\right)=b_{30}=d_{42}$,
${\rm v}_{P_{43}}\left(z\mathcal{O}_{K}\right)=b_{1}=d_{43}$, ${\rm v}_{\overline{P}_{43}}\left(z\mathcal{O}_{K}\right)=b_{16}=d_{43}$, ${\rm v}_{P_{44}}\left(z\mathcal{O}_{K}\right)=b_{13}+b_{51}=d_{44}$, ${\rm v}_{\overline{P}_{44}}\left(z\mathcal{O}_{K}\right)=b_{12}=d_{44}$, ${\rm v}_{P_{45}}\left(z\mathcal{O}_{K}\right)=b_{51}=d_{45}$,
${\rm v}_{\overline{P}_{45}}\left(z\mathcal{O}_{K}\right)=b_{18}+b_{40}=d_{45}$, ${\rm v}_{P_{46}}\left(z\mathcal{O}_{K}\right)=b_{7}+b_{10}+b_{32}+b_{33}+b_{35}+b_{49}+b_{52}=d_{46}$, ${\rm v}_{\overline{P}_{46}}\left(z\mathcal{O}_{K}\right)=b_{2}+b_{9}+b_{42}=d_{46}$,
${\rm v}_{P_{47}}\left(z\mathcal{O}_{K}\right)=b_{6}+b_{50}+b_{54}=d_{47}$, ${\rm v}_{\overline{P}_{47}}\left(z\mathcal{O}_{K}\right)=b_{43}+2b_{56}=d_{47}$, ${\rm v}_{P_{48}}\left(z\mathcal{O}_{K}\right)=b_{9}=d_{48}$, ${\rm v}_{\overline{P}_{48}}\left(z\mathcal{O}_{K}\right)=b_{5}+b_{12}=d_{48}$,
${\rm v}_{P_{49}}\left(z\mathcal{O}_{K}\right)=b_{27}+b_{41}=d_{49}$, ${\rm v}_{\overline{P}_{49}}\left(z\mathcal{O}_{K}\right)=b_{14}+b_{40}+b_{48}=d_{49}$, ${\rm v}_{P_{50}}\left(z\mathcal{O}_{K}\right)=b_{57}=d_{50}$, ${\rm v}_{\overline{P}_{50}}\left(z\mathcal{O}_{K}\right)=b_{24}+b_{49}=d_{50}$,
${\rm v}_{P_{51}}\left(z\mathcal{O}_{K}\right)=b_{5}+b_{22}=d_{51}$, ${\rm v}_{\overline{P}_{51}}\left(z\mathcal{O}_{K}\right)=b_{41}=d_{51}$, ${\rm v}_{P_{52}}\left(z\mathcal{O}_{K}\right)=b_{3}+b_{17}+b_{21}+b_{32}+b_{48}=d_{52}$, ${\rm v}_{\overline{P}_{52}}\left(z\mathcal{O}_{K}\right)=b_{8}+b_{10}=d_{52}$,
${\rm v}_{P_{53}}\left(z\mathcal{O}_{K}\right)=b_{3}+b_{11}=d_{53}$, ${\rm v}_{\overline{P}_{53}}\left(z\mathcal{O}_{K}\right)=b_{1}+b_{14}+b_{51}+b_{57}=d_{53}$, ${\rm v}_{P_{54}}\left(z\mathcal{O}_{K}\right)=b_{8}+b_{39}=d_{54}$, ${\rm v}_{\overline{P}_{54}}\left(z\mathcal{O}_{K}\right)=b_{25}=d_{54}$,
${\rm v}_{P_{55}}\left(z\mathcal{O}_{K}\right)=b_{15}+b_{28}=d_{55}$, ${\rm v}_{\overline{P}_{55}}\left(z\mathcal{O}_{K}\right)=b_{10}+b_{35}=d_{55}$, ${\rm v}_{P_{56}}\left(z\mathcal{O}_{K}\right)=b_{37}=d_{56}$, ${\rm v}_{\overline{P}_{56}}\left(z\mathcal{O}_{K}\right)=b_{26}+b_{47}=d_{56}$.